OpenStudy (anonymous):

my homework has a lot of equations 62 @zaibali.qasmi

OpenStudy (anonymous):

ask here...

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

X=6 ,X=-9 FOR x^2+3x-54=0

OpenStudy (anonymous):

i don't fully understand the question

OpenStudy (anonymous):

for x^2+4x+c=0 option D...c=-7

OpenStudy (anonymous):

9x^2=-4x+16 option C..

OpenStudy (anonymous):

ok thank you @zaibali.qasmi for all your help

OpenStudy (anonymous):

u welcome..